IN TRIANGLE ,ABC  the value ogf the following equation … COS( B+2C+3A)+COS(A-B) is 
 

Grade:11

1 Answers

Vijay Mukati
7 years ago
Dear Student,

Here you can write, B + 2C + 3A = pi + C + 2A = K(let),
so cos(K) = -cos(2A+C),
Therefore, now given equation becomes, cos(A-B) – cos (2A+C) now from here you can simplify using applying the formula of CosC – CosD.

Thanks, Vj